# Temporal-Spectral-Spatial Unified Remote Sensing Dense Prediction

### v1

@@ Abstract
The proliferation of diverse remote sensing data has spurred advancements in dense prediction tasks, yet significant challenges remain in handling data heterogeneity. Remote sensing imagery exhibits substantial variability across temporal, spectral, and spatial (TSS) dimensions, complicating unified data processing. Current deep learning models for dense prediction tasks, such as semantic segmentation and change detection, are typically tailored to specific input-output configurations. Consequently, variations in data dimensionality or task requirements often lead to significant performance degradation or model incompatibility, necessitating costly retraining or fine-tuning efforts for different application scenarios. This paper introduces the Temporal-Spectral-Spatial Unified Network (TSSUN), a novel architecture designed for unified representation and modeling of remote sensing data across diverse TSS characteristics and task types. TSSUN employs a Temporal-Spectral-Spatial Unified Strategy that leverages meta-information to decouple and standardize input representations from varied temporal, spectral, and spatial configurations, and similarly unifies output structures for different dense prediction tasks and class numbers. Furthermore, a Local-Global Window Attention mechanism is proposed to efficiently capture both local contextual details and global dependencies, enhancing the model’s adaptability and feature extraction capabilities. Extensive experiments on multiple datasets, encompassing tasks like building detection and land use/land cover classification with diverse TSS configurations, demonstrate that a single TSSUN model effectively adapts to heterogeneous inputs and unifies various dense prediction tasks. The proposed approach consistently achieves or surpasses state-of-the-art performance, highlighting its robustness and generalizability for complex remote sensing applications without requiring task-specific modifications.
@@
Remote Sensing, Dense Prediction, Unified Network, Change Detection, Semantic Segmentation, Deep learning
@@ I INTRODUCTION
With the continuous advancement of remote sensing technologies and the increasing diversity of data acquisition methods [1], the field of remote sensing has entered a phase of rapid development [2]. Massive and multi-source remote sensing data have been widely applied to various dense prediction tasks, playing a crucial role in applications such as urban expansion monitoring [3], land cover classification [4], disaster damage assessment [5], crop classification and yield estimation [6], and environmental pollution monitoring [7]. Remote sensing imagery exhibits high heterogeneity across three key dimensions—temporal, spectral, and spatial—posing significant challenges for unified processing due to variations in temporal length, number of spectral channels, and spatial resolution in practical applications [8].
@@
From the task perspective, dense prediction in remote sensing primarily involves three core categories: semantic segmentation, semantic change detection, and binary change detection. These tasks can be formally defined as follows: given a remote sensing image of shape $(T_{1},C_{1},H_{1},W_{1})$, the deep learning model is expected to produce a prediction of shape $(T_{2},C_{2},H_{2},W_{2})$, where $T_{1}$ and $T_{2}$ denote the temporal lengths of the input and output, $C_{1}$ is the number of input channels, $C_{2}$ is the number of output classes, and $(H_{1},W_{1})$ and $(H_{2},W_{2})$ denote the spatial dimensions of the input and output, respectively. Typically, $H_{1}=H_{2}$ and $W_{1}=W_{2}$. In semantic segmentation, the model performs multi-class land cover extraction from a single time-point image, corresponding to $T_{2}=1$ and $C_{2}\geq 2$; in semantic change detection, the model extracts land cover information at each time step to analyze semantic differences between any two time points, corresponding to $T_{2}=T_{1}$ and $C_{2}\geq 2$; in binary change detection, the model identifies whether changes occur between adjacent time points, corresponding to $T_{2}=T_{1}-1$ and $C_{2}=2$.
In recent years, deep learning methods, supported by the growing abundance of remote sensing data, have achieved notable success in dense prediction tasks, leading to the emergence of a wide range of high-performing models across all three task categories. However, existing models are typically designed for fixed input-output configurations, i.e., specific $(T_{1},C_{1},H_{1},W_{1})$ and $(T_{2},C_{2},H_{2},W_{2})$. Even slight variations in any of these dimensions can lead to significant performance degradation or complete incompatibility. In particular, changes in $T_{1}$, $C_{1}$, and $(H_{1},W_{1})$ reflect the diversity of remote sensing data along the Temporal- Spectral-Spatial (TSS) dimensions; variations in $T_{2}$ correspond to different task types, while differences in $C_{2}$ relate to varying classification requirements. Given that remote sensing applications often differ across these dimensions, models trained in one scenario typically struggle to generalize to others, requiring additional training or fine-tuning, which introduces considerable computational and time costs.
@@
Figure 1: Illustration of Temporal-Spectral-Spatial Unified Network. TSSUN is capable of handling input and output with arbitrary temporal, spectral, and spatial dimensions, supports dense prediction for any number of classes, and unifies semantic segmentation, binary change detection and semantic change detection tasks.
To address these issues, this study propose a Temporal-Spectral-Spatial Unified Network (TSSUN) that enables unified representation and modeling of remote sensing data across the temporal, spectral, and spatial dimensions, as shown in Figure 1. TSSUN is highly flexible, capable of handling any combination of $(T_{1},C_{1},H_{1},W_{1})$ inputs and $(T_{2},C_{2},H_{2},W_{2})$ outputs. It supports inputs with varying temporal lengths, spectral bands, and spatial resolutions, and unifies all three dense prediction tasks. Moreover, it can produce outputs with an arbitrary number of semantic classes and maintains strong performance across diverse configurations.
Specifically, we introduce the Temporal-Spectral-Spatial Unified Strategy (TSSUS) to decouple and unify the input and output representations across the TSS dimensions. At the input stage, TSSUS leverages inherent spectral and spatial characteristics to achieve unified encoding of remote sensing imagery. Along the spectral dimension, it exploits the continuity across spectral bands and integrates spectral metadata to accomplish spectral harmonization. In parallel, along the spatial dimension, it harnesses the similarity among nearby regions, supplemented with spatial metadata, to produce a unified representation of spatial scales. By preserving the time dimension at the input, TSSUS is able to independently extract features from remote sensing images across distinct temporal instances. At the encoder–decoder junction, TSSUS facilitates feature-level fusion along the temporal axis, effectively mitigating interference from irrelevant information while adapting the output temporal length to suit the dense prediction task at hand [9, 10]. At the output stage, TSSUS further capitalizes on the structured nature of remote sensing dense prediction tasks in both spectral and spatial dimensions. In the spectral dimension, by modeling inter-class correlations, it enables prediction over an arbitrary number of classes. Meanwhile, in the spatial dimension, by exploiting the spatial structural consistency among neighboring regions, it significantly enhances the model’s generalization capability when outputting data at varying spatial resolutions.
@@
To further improve the model’s ability to capture diverse TSS combinations, we design a Local-Global Window Attention (LGWA) mechanism. This module efficiently extracts local features using three overlapping window-based attention blocks with different shapes, followed by a global attention block that aggregates information at the global level. This design achieves a balance between computational efficiency and expressive power, enabling collaborative modeling of local and global features and significantly boosting the model’s performance in complex remote sensing tasks.
In summary, the main contributions of this work are as follows:
@@
    A Temporal-Spectral-Spatial Unified Network is proposed to enable unified modeling across temporal, spectral, and spatial dimensions for dense prediction in remote sensing. TSSUN supports arbitrary TSS input configurations and accommodates various tasks, including semantic segmentation, semantic change detection, and binary change detection with flexible output class settings.
@@
    A Local-Global Window Attention mechanism is designed to efficiently capture both local and global contextual features, improving performance across a wide range of remote sensing prediction tasks.
@@
    Extensive experiments are conducted on multiple datasets with diverse TSS configurations, including building detection and land use/land cover (LULC) classification. The results show that a single TSSUN model can adapt to heterogeneous inputs, unify multiple task types, and consistently achieve or exceed state-of-the-art (SOTA) performance.
@@ II Related Works
II-A Task Unification in Remote Sensing Dense Prediction
Dense prediction tasks in remote sensing, particularly semantic segmentation, binary change detection, and semantic change detection, form the cornerstone of many Earth observation applications. While often treated as distinct problems, they share fundamental objectives and challenges.
Semantic segmentation (SS) in remote sensing aims to assign a thematic label to each pixel in an image, thereby partitioning it into meaningful regions corresponding to different land cover or land use classes [11]. Early approaches relied on traditional machine learning algorithms such as Support Vector Machines (SVMs) and Random Forests (RFs) operating on handcrafted features [12]. The advent of deep learning, particularly Fully Convolutional Networks (FCNs) [13], revolutionized the field, enabling end-to-end feature learning and significantly improving segmentation accuracy. Architectures like U-Net [14], with its encoder-decoder structure and skip connections, became foundational, inspiring numerous variants such as UNet++ [15] and the DeepLab series [15, 16], which introduced atrous convolutions and spatial pyramid pooling to handle scale variations.
Binary change detection (BCD) focuses on identifying areas of significant difference between two co-registered remote sensing images acquired at different times [17]. Traditional BCD methods include image differencing, principal component analysis (PCA)-based techniques [18], and change vector analysis (CVA) [19]. Deep learning has brought substantial advancements, with Siamese networks being a popular choice for learning discriminative features from image pairs [20, 21]. These networks typically process the bi-temporal images through shared-weight encoders and then compare the resulting feature maps to identify changes. U-Net-like architectures have also been adapted for BCD, often by modifying the input layers to accept concatenated bi-temporal images or difference images, and training the network to output a binary change map [22].
Semantic change detection (SCD), also known as ”from-to” change detection, provides a more detailed analysis by not only identifying where changes have occurred but also classifying the type of land cover transition (e.g., forest to urban) [23]. This task is inherently more complex than BCD. A common traditional approach is post-classification comparison (PCC), where SS is performed independently on multi-temporal images, and the resulting maps are compared [24]. However, PCC suffers from error accumulation from the individual classification steps. More integrated methods aim to perform simultaneous segmentation and change identification. Deep learning models for SCD often employ dual-branch or multi-task learning frameworks, where separate decoders might be used for segmentation at each time step and for the final change map [25].
Critically analyzing these three tasks reveals profound interrelations and shared challenges. SCD inherently builds upon the principles of SS, as identifying ”from-to” transitions requires accurate semantic understanding of the landscape at different time points. BCD can be viewed as a simplified abstraction of SCD, where all types of semantic changes are aggregated into a binary ”change” or ”no change” outcome. Thus, accurate SS is a prerequisite for high-quality SCD, and insights from BCD can potentially guide the focus of SCD algorithms. The complementary nature of these tasks is evident: SS provides the static land cover context, BCD highlights dynamic regions, and SCD offers a comprehensive narrative of landscape transformation. Information from one task can, in principle, regularize or enhance the performance of others; for instance, knowing the semantic class of a pixel can refine change boundaries, or detecting a change can prompt a more detailed semantic analysis. Despite these intrinsic relationships and the potential for synergistic modeling, there is currently a significant lack of models or frameworks that can effectively unify semantic segmentation, binary change detection, and semantic change detection within a single, cohesive architecture.
II-B Data Dimensionality in Remote Sensing Dense Prediction
@@
Remote sensing datasets utilized for dense prediction tasks are characterized by extensive variability across temporal, spectral, and spatial dimensions. This heterogeneity stems from the diverse array of satellite and airborne sensors, each with unique acquisition parameters, mission objectives, and coverage patterns, tailored for different application scenarios and geographical regions [26]. Such variability presents a formidable challenge for developing universally applicable and robust dense prediction models.
@@
The temporal dimension in remote sensing data exhibits significant disparities. Revisit frequency, a critical factor for monitoring dynamic phenomena, ranges from multiple observations per day with sensors like MODIS [27], to several days (e.g., Sentinel-1 and Sentinel-2, with 5-12 day repeat cycles depending on latitude and constellation status [28, 29]), to 16 days for Landsat missions [30]. Consequently, the temporal length of image sequences available for analysis can vary from bi-temporal pairs, commonly used in BCD and SCD [[20], [23]], to dense time series comprising hundreds of observations, which are invaluable for applications like agricultural monitoring [31]. Models trained on data with a specific temporal frequency or sequence length may not generalize well to datasets with different temporal characteristics without substantial retraining or adaptation.
@@
Spectral dimensionality is another source of major variation. The number of spectral bands can range from a single panchromatic band to a few multispectral bands (e.g., 4-bands in NAIP imagery, 13 bands in Sentinel-2 MSI [29]), to hundreds of narrow, contiguous bands in hyperspectral sensors like AVIRIS [32] or the upcoming EnMAP mission [33]. Each sensor captures information from different portions of the electromagnetic spectrum, with varying band central wavelengths and bandwidths. This spectral diversity allows for the discrimination of different materials and land cover types based on their unique spectral signatures. However, it also means that models developed for one sensor may not be directly applicable to data from another sensor without strategies to handle the differing channel counts and spectral information content [34].
Spatial characteristics, primarily ground sampling distance (GSD) or resolution, also vary widely. Data can range from very high resolution (VHR), with GSDs of less than 1 meter, to high resolution (HR, 1-10 meters, e.g., Sentinel-2, PlanetScope), to lower resolution. VHR imagery is suitable for detailed urban mapping or object detection [35], while medium to low-resolution data is often used for regional or global land cover mapping and monitoring [36]. The spatial extent of individual scenes also differs, impacting the scale of analysis. Models trained on imagery of a particular GSD often struggle when applied to data with significantly different spatial resolutions, as object appearance and contextual cues change drastically.
@@
The inherent heterogeneity in temporal frequency, spectral composition, and spatial resolution across remote sensing datasets thus poses a substantial hurdle for developing universally applicable dense prediction models. Consequently, there is a pressing research gap concerning the development of adaptive model architectures or unified data processing strategies that can effectively ingest and interpret such diverse time-spectrum-space data formats at the input level for comprehensive and robust dense prediction.

### v2

# Spatial-Temporal-Spectral Unified Modeling for Remote Sensing Dense Prediction

@@ Abstract
The proliferation of multi-source remote sensing data has propelled the development of deep learning for dense prediction, yet significant challenges in data and task unification persist. Current deep learning architectures for remote sensing are fundamentally rigid. They are engineered for fixed input-output configurations, restricting their adaptability to the heterogeneous spatial, temporal, and spectral dimensions inherent in real-world data. Furthermore, these models neglect the intrinsic correlations among semantic segmentation, binary change detection, and semantic change detection, necessitating the development of distinct models or task-specific decoders. This paradigm is also constrained to a predefined set of output semantic classes, where any change to the classes requires costly retraining. To overcome these limitations, we introduce the Spatial-Temporal-Spectral Unified Network (STSUN) for unified modeling. STSUN can adapt to input and output data with arbitrary spatial sizes, temporal lengths, and spectral bands by leveraging their metadata for a unified representation. Moreover, STSUN unifies disparate dense prediction tasks within a single architecture by conditioning the model on trainable task embeddings. Similarly, STSUN facilitates flexible prediction across any set of semantic categories by integrating trainable category embeddings as metadata. Extensive experiments on multiple datasets with diverse STS configurations in multiple scenarios demonstrate that a single STSUN model effectively adapts to heterogeneous inputs and outputs, unifying various dense prediction tasks and diverse semantic class predictions. The proposed approach consistently achieves state-of-the-art performance, highlighting its robustness and generalizability for complex remote sensing applications.
@@
Remote Sensing, Dense Prediction, Data Unification, Task Unification, Change Detection, Semantic Segmentation, Deep learning
@@ I INTRODUCTION
With the continuous advancement of remote sensing technologies and the increasing diversity of data acquisition methods [1], the field of remote sensing has entered a phase of rapid development [2]. Massive and multi-source remote sensing data have been widely applied to various dense prediction tasks, playing a crucial role in applications such as urban expansion monitoring [3], land cover classification [4], disaster damage assessment [5], crop classification and yield estimation [6], and environmental pollution monitoring [7]. Remote sensing imagery exhibits high heterogeneity across three key dimensions including spatial, temporal and spectral dimensions, posing significant challenges for unified processing due to variations in image size, temporal length and spectral bands in practical applications [8].
@@
Remote sensing dense prediction refers to tasks where the goal is to produce a pixel-wise prediction, assigning a specific label to every pixel in the input satellite imagery [9]. From the task perspective, dense prediction in remote sensing primarily involves three core categories: semantic segmentation, binary change detection, and semantic change detection. These tasks can be formally defined as follows: given a remote sensing image time series of shape $(T_{1},C_{1},H_{1},W_{1})$, the deep learning model is expected to produce a prediction of shape $(T_{2},C_{2},H_{2},W_{2})$, where $T_{1}$ and $T_{2}$ denote the temporal lengths of the input and output, $C_{1}$ is the number of input channels, $C_{2}$ is the number of output classes, and $(H_{1},W_{1})$ and $(H_{2},W_{2})$ denote the image sizes of the input and output, respectively. Typically, $H_{1}=H_{2}$ and $W_{1}=W_{2}$. In semantic segmentation, the model aims to classify land cover types at a per-pixel level [9], corresponding to $T_{2}=1$ and $C_{2}\geq 2$; in binary change detection, the model identifies whether changes occur between adjacent time points [10], corresponding to $T_{2}=T_{1}-1$ and $C_{2}=2$; in semantic change detection, the model extracts land cover information at each time step to analyze semantic differences between any two time points [11], corresponding to $T_{2}=T_{1}$ and $C_{2}\geq 2$. Therefore, the temporal dimension $T_{1},T_{2}$ is strongly correlated with the type of dense prediction tasks, and the output spectral dimension $C_{2}$ is strongly correlated with the set of predicted semantic categories.
@@
Figure 1: Illustration of Spatial-Temporal-Spectral Unified Network. STSUN is capable of handling input and output with arbitrary spatial, temporal and spectral dimension configurations, unifies semantic segmentation, binary change detection and semantic change detection tasks with flexible category set.
While deep learning methods have achieved significant progress in remote sensing dense prediction, yielding high-performing models across semantic segmentation, binary change detection, and semantic change detection tasks [12, 13, 14, 15], current architectures exhibit several limitations. 1. Fixed Configurations: These models are typically designed for fixed input-output configurations, defined by specific image sizes $((H1,W1),(H2,W2))$, temporal lengths $(T1,T2)$ and spectral bands $(C_{1},C_{2})$. This rigidity restricts their adaptability to the diverse and heterogeneous remote sensing datasets encountered in practical applications, such as diverse satellite sensor data with varying spectral bands or multi-temporal imagery sequences with varying temporal lengths. 2. Fixed Task: These models neglect the inherent correlations among semantic segmentation, binary change detection, and semantic change detection tasks. This oversight often mandates the development of distinct models or dedicated task-specific decoders for each task, hindering unified processing pipelines. 3. Fixed Category Set: Existing approaches are generally constrained to dense prediction with a fixed set of output categories. Consequently, any alteration to the target dense prediction schema, even a slight change, often result in substantial performance degradation or complete incompatibility, necessitating extensive retraining or fine-tuning, which incurs considerable computational and temporal overhead.
To overcome these limitations, this study introduces the Spatial-Temporal-Spectral Unified Network (STSUN), as illustrated in Figure 1. STSUN offers a novel framework for unified representation and modeling of remote sensing data across diverse spatial, temporal and spectral dimensions, solving the above issues in the following ways. 1. Unified Representation: STSUN effectively addresses the variability in data characteristics by leveraging inherent metadata from each dimension to achieve a unified representation. This capability allows STSUN to seamlessly adapt to input and output data with arbitrary image sizes, temporal lengths, and spectral bands, thereby leveraging a wide range of data to build high-performance unified models for dense prediction. 2. Unified Tasks: STSUN unifies semantic segmentation, binary change detection, and semantic change detection tasks within a single framework, which is accomplished by using a predefined and trainable task embedding set. By incorporating a selected task embedding and output temporal length into the output temporal dimension as metadata, the model is explicitly instructed to execute on the exact dense prediction task and the necessary output temporal length. This approach eliminates the need for separate models or task-specific decoders, enabling effective multi-task joint modeling and leveraging extensive data from multiple dense prediction tasks to improve model performance. 3. Flexible Category Set: STSUN facilitates flexible prediction across an arbitrary set of semantic categories, which is accomplished by using a predefined and trainable category embedding set. The selected category subset is integrated into the output channel dimension as metadata, which guides the model to identify the specific classes needed for a given prediction. This enables the model to adapt to multiple subsets of prediction categories without fine-tuning, applying to diverse remote sensing scenarios with various semantic categories.
@@
To further improve the model’s ability to capture diverse STS combinations, we design a Local-Global Window Attention (LGWA) mechanism. This module efficiently extracts local features using three overlapping window-based attention blocks with different shapes, followed by a global attention block that aggregates information at the global level. This design achieves a balance between computational efficiency and expressive power, enabling collaborative modeling of local and global features and boosting the model’s performance in complex remote sensing tasks.
The main contributions of this work are as follows:
@@
    We propose the Spatial-Temporal-Spectral Unified Network, which accommodates arbitrary input and output configurations across spatial, temporal, and spectral dimensions for the first time. This addresses the prevalent issue of model rigidity in handling heterogeneous remote sensing data.
@@
    We introduce a unified task execution method that leverages trainable task embeddings to perform semantic segmentation, binary change detection, and semantic change detection within a single architecture, thereby obviating the need for separate models or task-specific heads.
@@
    We design a flexible category prediction method, which utilizes trainable semantic embeddings to enable the model to perform dense prediction for any specified set of output classes, removing the constraint of a fixed category schema and the associated costs of retraining.
  4. 4.
    We develop a Local-Global Window Attention mechanism that efficiently captures both local and global contextual features. This design enhances the model’s feature extraction capabilities and improves performance across a wide range of remote sensing prediction tasks.
@@ II Related Works
II-A Data Dimensionality in Remote Sensing Dense Prediction
@@
Remote sensing datasets utilized for dense prediction tasks are characterized by extensive variability across spatial, temporal and spectral dimensions. This heterogeneity stems from the diverse array of satellite and airborne sensors, each with unique acquisition parameters, mission objectives, and coverage patterns, tailored for different application scenarios and geographical regions [16]. Such variability presents a formidable challenge for developing universally applicable and robust dense prediction models.
Remote sensing imagery exhibits substantial spatial heterogeneity stemming from two key properties: ground sampling distance (GSD) and geospatial extent. The GSD is sensor-dependent, ranging from sub-meter resolutions ideal for detailed urban analysis [17] to coarser resolutions suited for regional land cover mapping [18]. The geospatial extent, meanwhile, is determined by the specific application, leading to variations in the captured area. Together, these properties dictate not only the pixel dimensions of an image but also the representation of its content. This variability poses a critical generalization challenge for deep learning models. A model trained for a specific resolution and image size typically suffers a significant performance degradation when applied to data with different spatial characteristics. This failure arises from two coupled effects: variations in GSD alter the perceived scale and texture of ground objects, while changes in geospatial extent modify the available contextual information.
@@
The temporal dimension in remote sensing data exhibits significant disparities. Revisit frequency, a critical factor for monitoring dynamic phenomena, ranges from multiple observations per day with sensors like MODIS [19], to several days (e.g., Sentinel-1 and Sentinel-2, with 5-12 day repeat cycles depending on latitude and constellation status [20, 21]), to 16 days for Landsat missions [22]. Consequently, the temporal length of image sequences available for analysis can vary from bi-temporal pairs, commonly used in building change detection [23, 24], to dense time series comprising hundreds of observations, which are invaluable for applications like agricultural monitoring [25], vegetation forecasting [26] and moving object detection [27]. Models trained on data with a specific temporal sequence length may not apply to datasets with different temporal characteristics without substantial retraining and adaptation.
@@
Spectral dimensionality is another source of major variation. The number of spectral bands can range from a single panchromatic band to a few multispectral bands (e.g., 4-bands in NAIP imagery, 13 bands in Sentinel-2 MSI [21]), to hundreds of narrow, contiguous bands in hyperspectral sensors like AVIRIS [28] or the upcoming EnMAP mission [29]. Each sensor captures information from different portions of the electromagnetic spectrum, with varying band central wavelengths and bandwidths. This spectral diversity allows for the discrimination of different materials and land cover types based on their unique spectral signatures [30]. However, it also means that models developed for one sensor may not be directly applicable to data from another sensor without strategies to handle the differing channel counts and spectral information content [31]. A series of models have been proposed to use hypernetworks to unify the inputs with different spectral bands, but ignored the unification of the output spectral bands [32, 33].
@@
The inherent heterogeneity in image size, temporal frequency, spectral composition across remote sensing datasets poses a substantial hurdle for developing universally applicable dense prediction models. Consequently, there is a pressing research gap concerning the development of adaptive model architectures or unified data processing strategies that can effectively ingest and interpret such diverse time-spectrum-space data formats at the input and output level for comprehensive and robust dense prediction.
II-B Task Unification in Remote Sensing Dense Prediction
Dense prediction in remote sensing encompasses a range of pixel-level interpretation tasks that are crucial for remote sensing applications such as environmental monitoring, urban planning, and disaster assessment [9]. Among these, three tasks are fundamental: semantic segmentation, binary change detection, and semantic change detection. Semantic segmentation aims to assign a specific class label to every pixel in the satellite images, producing a detailed land cover map [34]. Binary change detection, conversely, utilizes multi-temporal images to identify pixels where any form of change has occurred, outputting a binary map of change, no change [10]. Bridging these two is semantic change detection, which not only detects changes across multi-temporal images but also identifies the ”from-to” nature of the change, such as a ’forest’ pixel becoming a ’building’ pixel [35].
Given the significant conceptual overlap and inherent correlations among these tasks, a growing trend in the community is task unification through multi-task learning frameworks. Developing unified models to handle multiple dense prediction tasks simultaneously can lead to improved performance and efficiency [35, 13], which leverage shared representations to allow complementary information from one task to benefit others. A model trained jointly for multiple tasks can learn more robust feature extractors than a model trained on either task alone [36]. For instance, SFCCD features task-specific branches for building semantic segmentation and change detection [35]. It leverages paired data from both tasks for training, resulting in superior change detection performance compared to training the change detection branch solely with change detection data. Similarly, FCCDN incorporates multi-task branches for semantic segmentation and change detection, and it utilizes the weakly supervised results from the semantic segmentation branch to enhance change detection effectiveness [13]. The core benefit of task unification is the potential to create more powerful and generalizable models while reducing the need to develop and deploy multiple specialized networks.
Despite their promise, existing multi-task unification efforts suffer from critical limitations that hinder their scalability and effectiveness. The first issue is a pervasive dependence on paired data. Current models typically require that the training datasets are fully annotated for all constituent tasks, which means the exact same set of images must possess corresponding pixel-level labels for semantic segmentation, binary change, and semantic change. This stringent requirement drastically limits the pool of usable training data, as such comprehensively annotated, multi-task datasets are exceedingly rare and expensive to create. The second issue is the common architectural choice of using task-specific decoders. Many unified models employ a shared encoder to extract features but diverge into separate, specialized decoder heads for each task. This design can create information bottlenecks and prevent the model from fully exploiting the synergies and complementarities among the tasks at the deepest levels of feature decoding and synthesis.
Consequently, there is an absence of a task unified dense prediction model that can effectively unify multiple core tasks through a joint modeling paradigm that operates on non-paired data.
II-C Flexible Class Sets in Remote Sensing Dense Prediction
The objective of most remote sensing dense prediction tasks is to categorize ground objects according to a single predefined semantic class set. However, the definition of this class set is not universal, which is highly dependent on the specific application and the characteristics of the geographic scene. This variability is evident across different tasks and scenarios. For example, a simple building extraction task may only require a binary class set of building, background, while binary change detection operates on change, no change. In contrast, a standard Land Use/Land Cover (LULC) classification task might involve a more complex set, such as water, forest, cropland, urban, barren.
This variability is further compounded by scene-driven factors. The LULC class schema for a dense urban environment might need to include specific categories like commercial building, residential building, road, playground, vehicle, which would be irrelevant in a forest monitoring application. The latter might instead require a fine-grained class set like coniferous forest, bamboo forest, tea plantation, shrubland. While existing deep learning models have achieved impressive performance, they are almost universally designed for a specific and fixed class schema [15, 12, 14].
The primary limitation of these models is their inherent inflexibility with respect to the semantic class set. They are designed, trained, and optimized for a fixed vocabulary of predefined classes. The architecture of the model, particularly the final classification layer, is hard-wired to the number and identity of these classes. As a result, even a minor modification to the class set would render the pretrained model unusable for the new task, such as adding a ’wetland’ category to a LULC model or splitting the ’building’ class into ’residential’ and ’commercial’. Adapting the model requires retraining or fine-tuning on a new dataset that includes the modified class set.
Therefore, there lacks a unified remote sensing dense prediction model capable of flexibly adapting to diverse class sets across different tasks and scenes without necessitating costly and time-consuming retraining.